OpenStudy (butterflydreamer):

help please :) During a fishing trip Alex notices that the height h of the tide (in metres) is given by h=2 + 0.5cos(pi/4*t) where t is measured in hours from the start of the trip. Find the rate of change of the height 2 hours after the start of the trip.

OpenStudy (butterflydreamer):

do you just derive h and then sub in t = 2??

OpenStudy (butterflydreamer):

if you derive \[h = 2 + 0.5 \cos (\frac{ \pi }{ 4 } t)\] you get : \[h' = -\frac{\pi}{8} \sin( \frac{ \pi }{ 4 }t)\]

OpenStudy (kky10997):

now but t =2hr

OpenStudy (butterflydreamer):

so.. \[h' = \frac{ -\pi }{ 8 } \sin(\frac{ \pi }{ 2 }) = -\frac{ \pi }{ 8 }\] ?

OpenStudy (kky10997):

yeah

OpenStudy (butterflydreamer):

oh thank you (:

OpenStudy (kky10997):

:)

OpenStudy (butterflydreamer):

i just checked and it says that it is incorrect :/

rishavraj (rishavraj):

is it \[-0.5\frac{ \pi }{ 4 }\]

OpenStudy (butterflydreamer):

nope

rishavraj (rishavraj):

hold on wht is it then ??

rishavraj (rishavraj):

\[-\frac{ \pi }{ 8 }\]

OpenStudy (butterflydreamer):

.... no. -0.5(pi/4) is the same as -pi/8

rishavraj (rishavraj):

lol yeah bt whts the answer given??

OpenStudy (butterflydreamer):

i don't know what it is LOL but it comes up as being incorrect :/

rishavraj (rishavraj):

hmmmm @butterflydreamer i guess the differentiation is correct though

OpenStudy (butterflydreamer):

yeah. it's strange... not sure why the answer is wrong.. unless we're interpreting the question wrong